# NeRF-Aug: Data Augmentation for Robotics with Neural Radiance Fields

### v1

@@ V-C Real World Tasks
TABLE II: Time To Create New Data in Hours ($\downarrow$)
Grasping Pick and Place Drop Cup Pick Up
Methods Screwdriver Hammer Pliers Bottle Remote Eraser Spatula Green Plate Book Pan Box
Ours (first novel object) 5.2 5.3 5.0 5.3 5.3 5.7 5.3 4.3 4.2 4.6 4.0
Ours (subsequent novel objects) 2.1 2.2 1.9 2.5 2.5 2.9 2.5 2.0 1.9 2.3 1.8
GenAug 43.3 42.3 42.6 50.0 49.1 50.3 49.3 34.7 35.1 35.0 27.0
Instruct Pix-to-Pix 19.2 19.3 19.1 20.8 21.5 21.4 20.9 16.8 16.9 16.7 14.5
V-D Comparison to Baselines
As seen in Table I, our method consistently outperforms other methods on all four tasks. Fig. 7 and Fig. 5 demonstrate that by leveraging the photorealism of a NeRF blended with inpainted background, our method is able to create synthetic data with the same geometry and color as real-world data of a different object.
In contrast to our method, GenAug and Instruct Pix-to-Pix require intensive prompting to achieve a realistic looking image. We observe that the two methods are highly inconsistent from frame to frame because the random noise inputted into the diffusion model that leads to very different results each time a frame is rendered. The inconsistency can be detrimental to multi-camera systems where GenAug and Instruct Pix-to-Pix could render a completely different looking object for each camera image at the same time-step. Additionally, both GenAug and Instruct Pix-to-Pix seem to change the shading of the scene. Instruct Pix-to-Pix especially makes the background of the frames the same color as the object we prompted for.
Despite GenAug explicitly switching the texture of the object, many of the renders came out with an object that was the same color as the original object. We also notice that GenAug inpaints a new object with the exact geometry as the original object, which can be problematic because the new object is usually slightly bigger or smaller than the original one. This behavior also explains why GenAug works most effectively on picking and placing a remote which has the same dimensions and color as the original object (stapler). Finally, we argue that a major drawback of diffusion models in general is that they are not trained to render objects that appear very close to the camera. Gripper camera robotic systems, like our setup, will move the camera to only a few inches away from the object it is interacting with. Often only part of the object is in frame at that distance. We believe the camera was too close to the object for GenAug and Instruct Pix-to-pix to synthesize a recognizable bottle or remote in Fig. 7, especially in the later frames which look the most unrecognizable.
V-E Data augmentation speed
TABLE III: Breakdown of Hours Taken for Each Part Of
Our Method Averaged Per Novel Object
NeRF Segmentation NeRF
Dataset Training + Object Eraser Rendering
Grasp 0.69 3.06 1.37
Pick And Place 0.90 2.83 1.64
Drop 0.74 2.34 1.30
Pick Up 0.62 2.20 1.13
@@ V-E Data augmentation speed
V-F Is Ground-truth Object Position necessary?
TABLE IV: Ablation on Success Rates With Inferring
And Not Inferring Object Pose
Pick and Place
Methods Bottle Remote Eraser Spatula
Known Object Pose 100% 70% 40% 50%
Inferred Object Pose 70% 50% 30% 70%
Often, it is hard to precisely estimate the position of objects in robot demonstration data. As mentioned in Section IV-B, one way to estimate the object position is by recording the gripper position at the time the object was grasped as the starting object position. We conduct an ablation study on the pick and place task to show that estimating object pose by using a gripper is a reasonable approximation and does not immensely deteriorate the quality of the policy compared to using ground-truth object position. We report results in Table IV. Interestingly, the performance with the wooden spatula actually increases, likely because using the gripper to infer the object pose is more accurate than measuring it beforehand.
@@ VI Conclusion
We introduce NeRF-Aug, a data augmentation framework for gripper-camera robotic systems. Our method leverages the photorealism of NeRFs to replace training objects in expert demonstrations with novel objects. This allows us to create synthetic training data for novel objects that is virtually indistinguishable from data that would otherwise require a human to demonstrate. Through extensive quantitative and qualitative experiments on four real-world tasks with 11 different objects, we show that policies trained using NeRF-Aug data are consistently more successful at tasks involving novel objects than the baselines, while only requiring a fraction of the time to collect. Future research could explore other novel-view synthesis methods such as Gaussian Splatting and Plenoxels to generate similar augmentation frameworks.

### v2

@@ V-B Real World Tasks
TABLE II: Time To Create New Data in Minutes Including NeRF Model Training ($\downarrow$)
Remove Object Place in Chest Serve on Plate Cook with Pot Lift up High
Methods Blue Tongs Red Spoon Wood Stick Swiss Miss Box White Book Large Pan Sugar Box Red Mug Hammer
Ours (w/o inpainter) 12.77 15.0 12.8 13.2 15.2 12.8 11.4 5.87 5.36
Ours (w/ inpainter) 30.9 33.1 31.0 44.5 46.5 44.7 38.7 24.9 24.4
GenAug 556 548 555 572 566 480 419 116 117
Instruct Pix-to-Pix 73.5 73.6 70.4 75.4 76.5 64.7 55.9 15.7 15.8
Roboagent 1858 1835 1843 1913 1906 1633 1423 402 408
Dreambooth 751 738 746 777 781 660 526 160 156
V-C Comparison to Baselines
As seen in Table I, our method consistently outperforms other methods on all five tasks. Over the next best method, we achieve an average of 43% increase in the ”Remove Objects” task, a 5% increase in the ”Place In Chest” task, an 80% increase on the ”Cook with Pot” task, and a 50% increase on the ”Lift up High” task. Our method achieved a 85.6% success rate compared to the next best method (Dreambooth) with a 30% success rate.
Which methods synthesize object locations and sizes accurately? GenAug, Instruct Pix-to-Pix, and RoboAgent often output images of an object in either the wrong orientation or position. For example, as seen in fig.  2, we prompted the baselines to inpaint a ”red mug,” only for the output image to consist of a sideways view of a mug that is much larger than the actual mug. In contrast, our method was able to accurately synthesize the red mug from a top-down view. We argue that being able to render an object at unconventional viewpoints, such as when an object is very close to the camera or in a birds-eye-view, is a major reason why our method performs better than the baselines (such as being able to lift the mug 90% more often than any other method). Dreambooth seems more effective at synthesizing the correct object viewpoint, but will fail when the original object is significantly different in shape from the old object (see fig. 2 where Dreambooth fails to inpaint a swiss miss box and a book when the original object was a mustard bottle). A wide degree of viewpoints is important because images where the object is very close to the camera or from a birds-eye-view form the majority of an expert demonstration. Additionally, all of the baselines fail to inpaint the right size of the novel object.
Which method has higher image consistency? Moreover, we observe that aside from Dreambooth, the baselines are highly inconsistent in the color and presence of distractor objects from frame to frame. The inconsistency can be detrimental to multi-camera systems where these methods may render a different looking object for each camera image at the same time-step. See Fig. 6 for examples where the diffusion models inconsistently adds distractor objects in one image and no distractor objects later on. Roboagent especially seems to add many distractors to the background. On the other hand, NeRF-Aug keeps the same colors and does not add distractor objects.
How do the methods compare in synthesizing texture and color? We see that NeRF-Aug is able to accurately synthesize the exact color and texture of the novel object. Our most complicatedly textured objects were the Swiss Miss Box and Sugar Box, both of which NeRF-Aug was able to accurately synthesize down to the text on the box. In this area, GenAug fails by often incorporating the color of the original object into the modified image. Instruct Pix-to-Pix frequently makes the background of the frames the same color as the object we prompted for. Dreambooth seems to synthesize the texture correctly, but often fails to get the exact shade of the novel object.
V-D Data augmentation speed
TABLE III: Breakdown of Minutes Taken for Each Part of Our Method
Dataset Train Novel Object NeRF Segmentation Background NeRF Training + Inpainting Novel Object NeRF Rendering
Remove Object 3.40 4.75 13.38 9.72
Place In Chest 5.08 5.50 25.83 10.07
Serve On Plate 3.51 5.15 26.68 9.35
Lift Up High 5.07 1.45 17.61 2.07
Cook with Pot 3.03 4.47 22.92 8.36
@@ V-D Data augmentation speed
We also notice that the majority of the time taken was from the segmentation and background inpainter component (refer to Table III). This module only needs to run once per task as opposed to once per novel object because it deals with erasing the original object in the training videos which is shared across all novel objects. For this reason, we also calculated the time without the training and running of the inpainter, resulting in NeRF-Aug running 4.99 times faster than P2P. In Table II, we show the time taken including the segmentation, training the background NeRF model, and inpainting of the background NeRF (‘w/ inpainter’) and excluding this component (‘w/o inpainter’). We measure all times on a single Nvidia RTX A4000 graphics card.
@@ VI Conclusion
We introduce NeRF-Aug, a data augmentation framework for gripper-camera robotic systems. Our method leverages the photorealism of NeRFs to replace training objects in expert demonstrations with novel objects. This allows us to create synthetic training data for novel objects that is virtually indistinguishable from data that would otherwise require a human to demonstrate. Through extensive quantitative and qualitative experiments on 5 real-world tasks with 9 different objects, we show that policies trained using NeRF-Aug data are consistently more successful at tasks involving novel objects than the baselines, while only requiring a fraction of the time to augment. Future research could explore other novel-view synthesis methods such as Gaussian Splatting and Plenoxels to generate similar augmentation frameworks.